OpenStudy (anonymous):

Your cell phone company charges $50 per month for basic service, plus an additional $0.10 per text message. Let m represent the number of text messages and C represent the total monthly bill, write an equation to find the total monthly bill from your cell phone company.

OpenStudy (mckennarian):

C+(0.10x)=Monthly Bill

OpenStudy (anonymous):

which one @MckennaRian C = 0.10m - 50 C = 50m - 0.10 C = 0.10m + 50 C = 50m + 0.10